Total Manufacturing Cost, Income Statement, Unit Cost, and Selling Price
Two inventors, recently organized as Innovation, Inc., consult you regarding a planned new product.They have estimates of the costs of materials, labor, overhead, and other expenses for 2016 but need to know how much to charge for each unit to earn a profit in 2016 equal to 15% of their estimated total long-term investment of $480,000 (ignore income taxes).

Their plans indicate that each unit of the new product requires the following:

Direct Material 4 lb. of a material costing $6 per lb.
Direct labor 2 hrs. of a metal former's time at $13.20 per hr.
0.6 hr. of an assembler's time at $9.60 per hr.

Major items of production overhead would be annual rent of $55,752 for a factory building, $34,392 rent for machinery, and $26,040 of indirect material. Other production overhead is estimated to be $279,936. Selling expenses are an estimated 30% of total sales, and non-factory administrative expenses are 20% of total sales.


The consensus at Innovation is that during 2016 10,000 units of product should be produced for selling and another 2,000 units should be produced for the next year's beginning inventory. Also, an extra 3,000 pounds of material will be purchased as beginning inventory for the next year. Because of the nature of the manufacturing process, all units started must be completed, so work in process inventories are negligible.

Required
a. Incorporate the above data into a schedule of estimated total manufacturing costs and compute the unit production cost for 2016.

Do not use negative signs with any of your answers.

b. Prepare an estimated income statement that would provide the target amount of profit for 2016.

c. What unit sales price should Innovation charge for the new product?
Answer #1

Answer a.

Direct material:

Required for production during 2016 = 12,000 * 4lbs * $6 = $288,000

Required ending inventory = 3000 lbs * $6 = $18,000

Indirect material = $$26,040

Hence Purchase = $288,000 + $18,000 + $26040 = $332,040

Direct labor = No of production units * (2 hrs. of a metal former's time at $13.20 per hr.+ 0.6 hr. of an assembler's time at $9.60 per hr.)

= 12000 * (2 * 13.20 + 0.6 * 9.60)

= $385,920

Schedule of estimated total manufacturing costs is as follows:

Innovation, Inc Schedule of Estimated Cost of goods Manufactured For the year ended December 31, 2016 Cost of Goods Manufactured: Direct Material Issued/ used Beginning Direct material Invento Add, Purchases Direct Material available for issue/use Less, Indirect material Less, Ending Direct material Invento $0 $332,040 $332,040 $26,040 $18,000 $288,000 $385,920 Direct Labor Manufacturing overhead Rent - Factory buildirn Rent - Machine Indirect material Other production overhead $55,752 $34,392 $26,040 $279,936$396,120 $1,070,040 Cost of Goods Manufactured

Unit production cost for 2016 = Cost of goods manufactured / Number of units produced = 1070040 /12000 = $89.17

Answer b:

Target profit required for 2016 = 15% of their estimated total long-term investment of $480,000 = 480000 * 15% =$72,000

Let us assume sale value required = X

Selling expenses are an estimated 30% of total sales, and non-factory administrative expenses are 20% of total sales

Hence selling expense = 0.3X and non-factory administrative expenses = 0.2 X

Cost of goods sold as calculated below = $891,700

Hence:

X - 891700- 0.3X - 0.2X = 72000

=> X =$19,27,400

Innovation, Inc Cost of Goods Sold For the year ended December 31, 2016 Cost of Goods Sold: Direct Material Issued/ used Beginning Direct material Invento Add, Purchases Direct Material available for issue/use Less, Indirect material Less, Ending Direct material Invento $0 $332,040 $332,040 $26,040 $18,000 $288,000 $385,920 Direct Labor Manufacturing overhead Rent - Factory buildin Rent - Machine Indirect material Other production overhead $55,752 $34,392 $26,040 $279,936 $3 396,120 Cost of Goods Manufactured $1,070,040 $1,070,040 $891,700 Add, Beginning Finished Goods Invento Goods available for sale $0 Less, Ending Finished goods Invento Cost of Goods Sold: $178,340

Estimated income statement is as follows:

Innovation, Inc Statement of Income For the year ended December 31, 2016 Sale $1,927,400 Cost of Goods Sold: Direct Material Issued/ used Beginning Direct material Invento Add, Purchases Direct Material available for issue/use Less, Indirect material Less, Ending Direct material Invento $0 $332,040 $332,040 $26,040 $18,000 $288,000 $385,920 Direct Labor Manufacturing overhead Rent - Factory buildin Rent - Machine Indirect material Other production overhead $55,752 $34,392 $26,040 $279,936 $3 396,120 $1,070,040 Cost of Goods Manufactured Add, Beginning Finished Goods Invento Goods available for sale $0 $1,070,040 Less, Ending Finished goods Invento Cost of Goods Sold: Gross Income $178,340 $891,700 $1,035,700 $578,220 $385,480 $72,000 Selling expenses Non-factory administrative expenses Profit before Tax

Answer c:

Unit sales price should Innovation charge for the new product = Sale value / Estimated units sold

=1927400 /10000

= $192.74